# Turbulence in stratified rotating topographic wakes

### v1

@@ 3.3 Contribution of the centrifugal/inertial instability (CI)
Planetary rotation substantially changes the spatial organization of flow structures that are large enough to feel it ($Ro<O(10)$), and alters the dissipation in various ways, both direct and indirect. The direct effect of rotation can be sensed by the CI, whose intensity has a parametric dependence on rotation and is stronger at $Ro=O(1)$ (see Appendix C). The indirect effect, although still associated with CI, comes from the change of the mean/base flow through the modification of the coherent structures. For the latter, eddies that rotate in the same direction as the system rotation (cyclonic vortices, CVs) and those rotate in the opposite direction (anticyclonic vortices, AVs) behave differently. Among flows with $|\omega_{z}|=O(f)$, anticyclonic vortices and shear are often subject to CI, leading to an appreciable asymmetry between the two sides of the wake. We note that in the present wake, rotation is Southern Hemisphere and AVs are primarily on the right-hand-side of the flow ($y<0$ and $\omega_{z}>0$).
@@
where $\bm{\omega}=\nabla\times\bm{u}$ is the relative vorticity and $\tilde{b}=b+N^{2}z$ is the ‘total buoyancy’. PV is a useful diagnostic in large-scale flow analyses since it is conserved along isopycnal surfaces in the absence of friction or mixing. On the other hand, the sign of PV has implications for scales smaller than the balanced motions. Negative PV serves as an indicator of several hydrodynamic instabilities (Thomas et al., 2013), that are precursors to turbulence. The horizontal and vertical components of PV,
@@
Figure 7: Instantaneous dissipation rate $\varepsilon$ (colored) and its time-average $\langle{\varepsilon}\rangle$ (TKE dissipation, gray), probed on the horizontal line $z/h=0.25,x=1$. The time instances are the same as those in Fig. 6(a-d).
@@
Figure 5 shows PV on the horizontal plane $z/h=0.25$ at $Ro=0.75$ and different $Fr$. In the near wake of each case, entangled fine-scale structures of positive and negative PV can be seen, as a result of near wake turbulence. As the flow evolves into the intermediate wake, large patches of negative PV (indicated by blue color) can be seen on the anticyclonic side ($y<0$), with the magnitude decaying in the streamwise direction as the flow gradually adjusts to rotation, eventually reaching a near-zero PV, stable state with little turbulence. On the cyclonic side ($y>0$), strong large-scale coherent CVs can be found with positive PV, and they remain intact from destruction during their downstream propagation. This asymmetry between AVs and CVs is a characteristic feature of the SMS.
@@
At the $Fr\leq O(1)$ values of this case study, the flow exhibits a von Kármán shedding pattern in horizontal planes. However, there is a subtle difference in the vortex dynamics. At the lowest $Fr$ ($Fr=0.075$, Fig. 5(a)), dipoles are formed due to strong mutual interaction between the AVs and the CVs, where the CVs are systematically stronger and they attract AVs to the cyclonic side, leading to the veering of the wake. The increased horizontal vortex–vortex interaction that is closer to 2D dynamics is a consequence of strong stratification and limited vertical motions. This behavior is similar to the Bu25 case in Liu et al. (2024), where the vortices were tracked in time and dipole formation was statistically shown by the mean vortex trajectories and conditional vorticity distribution. At $Fr=0.15,0.30$ (Fig. 5(b-c)), the overall shedding pattern mimics that of a standard Kármán street while the CI of the AVs is more appreciable visually. At $Fr=0.40$ (Fig. 5(d)), the recirculation zone is significantly longer than that in other cases (which can also be seen in Fig. 2) and the vortex shedding pattern is less regular. The former is due to the fact that the hydraulic jet reaches a lower downward distance at lower $Fr$, and it interacts strongly with the separation (Chomaz et al., 1993). The vertical PV (${\it\Pi}_{v}$, not shown) is similar to the total PV, and the horizontal PV (${\it\Pi}_{h}$, not shown) did not show evidence of the symmetric instability.
@@
Figure 6 shows instantaneous $\varepsilon$ at $Fr=0.30$ and various $Ro$, where both the Kármán vortices and the shear layers present instabilities due to rotation. In all cases, small-scale turbulence structures can be seen in the near wake, whose spatial location coincides well with those wake eddies and the shear layer. These worm-like structures mimic the intense, randomly oriented vortex tubes in isotropic turbulence and are indicative of fully triggered turbulence as a consequence of the breakdown of the 3D instabilities and the establishment of a forward cascade. However, wake turbulence decays in the streamwise direction due to the stabilizing effect of strong stratification. Hence, these fine-scale structures don’t persist long after their generation and turbulence is localized to the near wake, which can also be seen in Fig. 5. The time-evolution of $\varepsilon$ and $\varepsilon_{\rho}$ for panels (a-d) can be found in the supplementary materials (movies). The spatial structure of $\varepsilon_{\rho}$ is similar to $\varepsilon$.
@@
Now we turn our attention to $y$-$z$ planes in the near wake ($x/D=1$), as shown in Fig. 6(e-g), in which the structures of CI-induced dissipation can be seen more clearly. For cases Fr030Ro1p5 and Fr030Ro3 in (b,c), rolled up dissipation structures can be seen at various heights on the anticyclonic (left) side, while the cyclonic shear layer is more stable (at this $x/D$ location). These dissipation structures are associated with the quasi-streamwise vortices during the growth period of the CI, and were also found in other rotating flows, e.g., in Kloosterziel et al. (2007); Arobone and Sarkar (2012); Carnevale et al. (2013). A linear analysis of the growth of the streamwise/azimuthal vorticity during the CI is provided in Appendix C, which demonstrates its relevance to the diagnosis of CI and the relation between the growth rates and respective instability criteria.
Figure 8: Volume-integrated dissipation, $\int_{\mathcal{V}}\langle\tilde{\varepsilon}\rangle\,{\rm d}V$, as a function of (a) $Ro$ and (b) $FrRo$. Here the integration domain is $\mathcal{V}=[-D,4D]\times[-2D,2D]\times[0,2h]$. The vertical bars denote one standard variance above and below the averages. The green and red shades in (b) mark the overlap of the present parameters and those in Perfect et al. (2020) and Srinivasan et al. (2021), respectively.
@@ 4 Parametric dependence of dissipation on stratification and rotation
is employed. The size of the integration subdomain is $\mathcal{V}=[-D,4D]\times[-2D,2D]\times[0,2h]$ that encloses the turbulent near wake. Here $\langle\tilde{\varepsilon}\rangle$ is the time-averaged (over $N_{t}\approx 30$ for 3D snapshots) dissipation rate of the total kinetic energy (instead of TKE), for better converged statistics and greater relevance to field measurements. For case Fr040, in which the recirculation bubble is longer and separation is later downstream, (13) is accordingly computed in an elongated domain $\mathcal{V}^{\prime}=[-D,8D]\times[-2D,2D]\times[0,2h]$ and no significant difference is found.
@@
Figure 8 shows the time- and volume-integrated dissipation rate $\mathcal{E}$ for each $Fr$-series as a function of $Ro$ in (a) and $FrRo$ in (b). It can be seen that the data break into two groups: (1) the lower-$Fr$ group, Fr007 and Fr015, and (2) the higher-$Fr$ group, Fr030 and Fr040.
@@
In the first group with very strong stratification, the effect of rotation appears to be solely stabilizing. The dissipation increases as $Ro$ increases, and eventually saturates at high $Ro$ or weak rotation. The dissipation in Fr007 cases is consistently lower than that in Fr015 cases, even though the $Re$ is twice higher in the Fr007 cases. The trends of monotonic decrease of dissipation with decreasing $Ro$ and $Fr$ are qualitatively consistent with the findings and the parameter ranges in previous ROMS studies of Perfect et al. (2020) and Srinivasan et al. (2021). A difference is that the weak rotation (higher $Ro$) saturation was not observed in the previous work since it concerned the strong rotation/large topography regime of $Fr,Ro\leq O(0.1)$. In Fig. 8(b), the green and red shaded regions mark the upper end of the parameter combinations of $FrRo$ in Perfect et al. (2020) and Srinivasan et al. (2021).
@@
In the second group where stratification is not too strong, the effect of rotation is non-monotonic – there is an intermediate value of $Ro$ where the volume-integrated dissipation peaks. The intermediate value is $Ro=0.40$ for the Fr040 cases and $Ro=0.75$ for the Fr030 cases. Both peaks fall within $Ro=O(0.5-1)$, corresponding to SMS topographies or eddies. The dissipation maxima at the SMS and the associated most destabilizing rotation have not been well explored in the previous parameterizations of topographic wakes and highlight the significance of CI and its $Ro$-dependence.
@@
Similar non-monotonic rotation dependence was seen in the rotating horizontal shear layer with and without vertical stratification (Yanase et al., 1993; Arobone and Sarkar, 2012), and is characteristic of CI. The numerical stability analysis of Yanase et al. (1993) and Arobone and Sarkar (2012) revealed that, with no rotation, the three-dimensionally most unstable mode is the 2D KH mode ($k_{z}=0$), which is still the case when there is vertical stratification (Arobone and Sarkar, 2012). As the authors found, when there is weak rotation ($Ro=O(10)$), the most unstable mode is still the KH mode, but a nearly streamwise invariant inertial mode starts to emerge. As the rotation rate increases and $Ro$ reaches $O(1)$, the growth rate of the inertial instability overtakes that of the KH mode. However, when rotation is further increased to $Ro\sim O(0.1)$, the inertial mode disappears. The existence of a most destabilizing rotation rate can be shown in a linear analysis of both parallel and circular flows in Appendix C, which explains the existence of an SMS dissipation peak in the Fr030 and Fr040 cases when stratification is not too strong.
As stratification becomes stronger, such as in cases Fr007 and Fr015, the vertical length scale decreases (see Fig. 2(a-d)) and the space for CI existence shrinks until it is eventually suppressed. Hence, the reason for rotation appearing to be solely stabilizing at $Fr=0.075,0.15$ is the absence of CI at strong stratification.
@@
The present parametric dependence has several implications. First, the inertial scaling $\varepsilon=C_{\varepsilon}U^{3}_{\infty}/D$, where $C_{\varepsilon}$ is a scaling coefficient, works reasonably well for the wake dissipation over a wide range in the parameter space of $(Fr,Ro)$ (the value of $FrRo$ varies by 2 orders of magnitude). It provides a means for contextualizing field measurements or modeling wake dissipation/mixing in regional or global climate models. On the other hand, the variability of $C_{\varepsilon}$ over the range of $Fr$ and $Ro$ investigated reflects the comprehensive effects of stratification and rotation. It is evident that turbulence in seamount or hill wakes at different levels of stratification depends differently on rotation, with a transition point between $Fr=0.15$ and $Fr=0.30$ due to the activation of the CI. That being said, $C_{\varepsilon}$ is likely not a simple function of $Fr$ and $Ro$.
@@
In previous numerical studies of topographic wakes (Perfect et al., 2018, 2020; Srinivasan et al., 2021), it was suggested that the vortex dynamics and turbulent dissipation might both be categorized by a single parameter, instead of $(Fr,Ro)$. For example, Perfect et al. (2018) varied both $Fr$ and $Ro$ in the range of $0.014\leq Fr^{*}\leq 0.14$ and $0.053\leq Ro^{*}\leq 0.21$, which is equivalent to $0.014\leq Fr\leq 0.14$ and $0.0265\leq Ro\leq 0.105$ after a conversion to our definition ($Fr=Fr^{*},Ro=Ro^{*}/2$). It was suggested that the vortex dynamics can be characterized as a function of the Burger number, $Bu=(Ro/Fr)^{2}=(Nh/f_{c}D)^{2}$, and the dissipation can further be parameterized as a function of a positive power of $FrRo$. Srinivasan et al. (2021) studied topographic wakes as a function of $Ro$ ($0.025\leq Ro\leq 1$) when $Fr=0.02$ is fixed, and found that the dissipation monotonically increases as a function of the Rossby number. For comparison, the same data in Fig. 8(a) is plotted in Fig. 8(b), as a function of $FrRo$. Although the data seems to collapse better, the division into two groups with different dependencies is still very clear. The parameters in Perfect et al. (2020) have $(FrRo)_{\rm max}\approx 0.015$ and those in Srinivasan et al. (2021) have $(FrRo)_{\rm max}=0.02$. Both overlap with the lower end of the present parameters, as shown by the green and red shades in Fig. 8(b).
Figure 9: Joint p.d.f (JPDF) of the total KE dissipation ($\tilde{\varepsilon}$, defined in (9)) and the local gradient Richardson number ($Ri_{g}$, defined in (10)). Contours enclose $85\%$ of the JPDF in each case. Plane location $z/h=0.25$. The $Fr$-dependent black dashed lines in each figure are given by $\tilde{\varepsilon}D/U_{\infty}^{3}=C_{R}Fr^{-1}Ri_{g}^{-1}$, with the same fitting constant $C_{R}=1/2500$.
@@
It was pointed out by Liu et al. (2024) that there is still vertical coupling of the vortex shedding at $Fr=0.15,\,Bu=\infty$ and hence the $Bu$-determination of coupling/decoupling is incomplete without an additional $Fr$-dependence – stronger stratification than $Fr=0.15$ is required to vertically decouple the vortex shedding in rotating and non-rotating Kármán wakes. A similar role of stratification is found in the dissipation dependence that the activation of the CI occurs between $Fr=0.15$ and $Fr=0.30$, which contributes greatly to dissipation and changes its $Ro$-dependence.

### v2

@@ 3.3 Contribution of the centrifugal/inertial instability (CI)
Planetary rotation substantially changes the spatial organization of flow structures that are large enough to feel it ($Ro<O(10)$), and alters the dissipation in various ways, both direct and indirect. The direct effect of rotation can be sensed by CI, whose intensity has a parametric dependence on rotation and is stronger at $Ro=O(1)$ (see Appendix C). The indirect effect, although still associated with CI, comes from the change of the mean/base flow through the modification of the coherent structures. For the latter, eddies that rotate in the same direction as the system rotation (cyclonic vortices, CVs) and those rotate in the opposite direction (anticyclonic vortices, AVs) behave differently. Among flows with $|\omega_{z}|=O(f)$, anticyclonic vortices and shear are often subject to CI, leading to an appreciable asymmetry between the two sides of the wake. We note that in the present wake, planetary rotation is clockwise as in the Southern Hemisphere and AVs are primarily on the right-hand side ($y<0$ and $\omega_{z}>0$).
@@
where $\bm{\omega}=\nabla\times\bm{u}$ is the relative vorticity and $\tilde{b}=b+N^{2}z$ is the ‘total buoyancy’. PV is a useful diagnostic in large-scale flow analyses since it is conserved along isopycnal surfaces in the absence of friction or mixing. On the other hand, the sign of PV has implications for scales smaller than the balanced motions. PV of opposite sign to $f_{c}$ serves as an indicator of several hydrodynamic instabilities (Thomas et al., 2013), that are precursors to turbulence. The horizontal and vertical components of PV,
@@
Figure 8: Instantaneous dissipation rate $\varepsilon$ (colored) and its time-average $\langle{\varepsilon}\rangle$ (TKE dissipation, gray), probed on the horizontal line $z/h=0.25,x/D=1$. The time instances are the same as those in Fig. 7(a-d).
@@
Figure 6 shows PV on the horizontal plane $z/h=0.25$ at $Ro=0.75$ and different $Fr$. In the near wake of each case, entangled fine-scale structures of positive and negative (normalized) PV can be seen, as a result of near wake turbulence. As the flow evolves into the intermediate wake, large patches of negative PV (indicated by blue color) can be seen on the anticyclonic side ($y<0$), with the magnitude decaying in the streamwise direction as the flow gradually adjusts to rotation, eventually reaching a near-zero PV, stable state with little turbulence. On the cyclonic side ($y>0$), strong large-scale coherent CVs can be found with positive PV, and they remain intact from destruction during their downstream propagation. This asymmetry between anticyclonic and cyclonic vorticity is a characteristic feature of submesoscale rotating flows with $Ro=O(1)$ and it will be analyzed quantitatively in section 4.
@@
At the $Fr\leq O(1)$ values of this case study, the flow exhibits a von Kármán shedding pattern in horizontal planes. Additionally, the case-dependent $Fr$ introduces a subtle difference in the vortex dynamics. At the lowest FrFr ($Fr=0.075$, Fig. 6(a)), dipoles are formed due to strong mutual interaction between the AVs and the CVs, where the CVs are systematically stronger and they attract AVs to the cyclonic side, leading to the veering of the wake. The increased horizontal vortex–vortex interaction that is closer to 2D dynamics is a consequence of strong stratification and limited vertical motions. This behavior is similar to the Bu25 case in Liu et al. (2024), where the vortices were tracked in time and dipole formation was statistically shown by the mean vortex trajectories and conditional vorticity distribution. At $Fr=0.15,0.30$ (Fig. 6(b-c)), the overall shedding pattern mimics that of a standard Kármán street while the CI of the AVs is more appreciable visually. At $Fr=0.40$ (Fig. 6(d)), the recirculation zone is significantly longer than that in other cases (which can also be seen in Fig. 3) and the vortex shedding pattern is less regular. The former is due to the fact that the hydraulic jet reaches a lower downward distance at higher $Fr$, and it interacts strongly with the separation (Chomaz et al., 1993). The vertical PV (${\it\Pi}_{v}$, not shown) is similar to the total PV, and the horizontal PV (${\it\Pi}_{h}$, not shown) did not show evidence of the symmetric instability.
@@
Figure 7 shows instantaneous $\varepsilon$ at $Fr=0.30$ and various $Ro$, where both the Kármán vortices and the shear layers present instabilities due to rotation. In all cases, small-scale turbulence structures can be seen in the near wake, whose spatial location coincides well with those wake eddies and the shear layer. These worm-like structures mimic the intense, randomly oriented vortex tubes in isotropic turbulence and are indicative of fully triggered turbulence as a consequence of the breakdown of the 3D instabilities and the establishment of a forward cascade. However, wake turbulence decays in the streamwise direction due to the stabilizing effect of strong stratification. Hence, these fine-scale structures do not persist long after their generation and turbulence is localized to the near wake, which can also be seen in Fig. 6. The time-evolution of $\varepsilon$ and $\varepsilon_{\rho}$ for panels (a-d) can be found in the supplementary materials (movies). The spatial structure of $\varepsilon_{\rho}$ is similar to $\varepsilon$.
@@
Now we turn our attention to $y$-$z$ planes in the near wake ($x/D=1$), as shown in Fig. 7(e-g), in which the structures of CI-induced dissipation can be seen more clearly. For cases Fr030Ro1p5 and Fr030Ro3 in (b,c), rolled up dissipation structures can be seen at various heights on the anticyclonic (left) side, while the cyclonic shear layer is more stable (at this $x/D$ location). These dissipation structures are associated with the quasi-streamwise vortices during the growth period of the CI, and were also found in other rotating flows, e.g., in Kloosterziel et al. (2007); Arobone and Sarkar (2012); Carnevale et al. (2013). A linear analysis of the growth of the streamwise/azimuthal vorticity during CI is provided in Appendix C, which demonstrates its relevance to the diagnosis of CI and the relation between the growth rates and respective instability criteria.
Figure 9: Volume-integrated dissipation, $\int_{\mathcal{V}}\langle\tilde{\varepsilon}\rangle\,{\rm d}V$, as a function of (a) $Ro$ and (b) FrRoFrRo. Here the integration domain is $\mathcal{V}=[-D,4D]\times[-2D,2D]\times[0,2h]$. The vertical bars denote one standard variance above and below the averages. The green and red shades in (b) mark the overlap of the present parameters and those in Perfect et al. (2020) and Srinivasan et al. (2021), respectively. (c-f): similar to (a), but the volume $\mathcal{V}_{\rm AC}$ contains only the cyclonic ($y>0$, marked with ‘C’) or the anticyclonic ($y<0$, marked with ‘A’) side.
@@ 4.1 Volumetric measures
is employed. The size of the integration subdomain is $\mathcal{V}=[-D,4D]\times[-2D,2D]\times[0,2h]$ that encloses the turbulent near wake. Here $\langle\tilde{\varepsilon}\rangle$ is the time-averaged (over $N_{t}\approx 30$ for 3D snapshots) dissipation rate of the total kinetic energy (instead of TKE), for better converged statistics and greater relevance to field measurements.
@@
Figure 9(a,b) show the time- and volume-integrated dissipation rate $\mathcal{E}$ for each $Fr$-series as a function of $Ro$ in (a) and $FrRo$ in (b). It can be seen that the data break into two groups: (1) the lower-$Fr$ group, Fr007 and Fr015, and (2) the higher-$Fr$ group, Fr030 and Fr040.
@@
In the first group with very strong stratification, the effect of rotation appears to be solely stabilizing. The dissipation increases as $Ro$ increases, and eventually saturates at high $Ro$ or weak rotation. The dissipation in Fr007 cases is consistently lower than that in Fr015 cases, even though the $Re$ is twice as high as in the Fr007 cases. The trends of monotonic decrease of dissipation with decreasing $Ro$ and $Fr$ are qualitatively consistent with the findings and the parameter ranges in previous ROMS studies of Perfect et al. (2020) and Srinivasan et al. (2021). A difference is that the weak rotation (higher $Ro$) saturation was not observed in the previous work since it concerned the strong rotation/large topography regime of $Fr,Ro\leq O(0.1)$. In Fig. 9(b), the green and red shaded regions mark the upper end of the parameter combinations of $FrRo$ in Perfect et al. (2020) and Srinivasan et al. (2021).
@@
In the second group where stratification is not too strong, the effect of rotation is non-monotonic – there is an intermediate value of $Ro$ where the volume-integrated dissipation peaks. The intermediate value is $Ro=0.40$ for the Fr040 cases and $Ro=0.75$ for the Fr030 cases. Both peaks fall within Ro=O(0.51)Ro=O(0.5-1), corresponding to SMS topographies or eddies. The dissipation maxima at the SMS and the associated most destabilizing rotation have not been well explored in the previous parameterizations of topographic wakes and highlight the significance of CI and its $Ro$-dependence.
@@
Similar non-monotonic rotation dependence was seen in the rotating horizontal shear layer with and without vertical stratification (Yanase et al., 1993; Arobone and Sarkar, 2012), and is characteristic of CI. The numerical stability analysis of Yanase et al. (1993) and Arobone and Sarkar (2012) revealed that, with no rotation, the three-dimensionally most unstable mode is the 2D KH mode ($k_{z}=0$), which is still the case when there is vertical stratification (Arobone and Sarkar, 2012). As the authors found, when there is weak rotation ($Ro=O(10)$), the most unstable mode is still the KH mode, but a nearly streamwise invariant inertial mode starts to emerge. As the rotation rate increases and $Ro$ reaches $O(1)$, the growth rate of the inertial instability overtakes that of the KH mode. However, when rotation is further increased to $Ro\sim O(0.1)$, the inertial mode disappears. The existence of a most destabilizing rotation rate can be shown in a linear analysis of both parallel and circular flows in Appendix C, which explains the existence of a dissipation peak at $Ro=O(0.5)$ in the Fr030 and Fr040 cases when stratification is not too strong. Cases Fr007 and Fr015 do not have this behavior.
Unlike studies focused solely on CI (Yanase et al., 1993; Kloosterziel et al., 2007; Arobone and Sarkar, 2012; Carnevale et al., 2013) where cyclonic and anticyclone shear/vortices are treated separately, the present wakes contain both, leading to anticipated asymmetry. Thus, the volumetric measure in Fig. 9(a) is repeated in (c-f) but with the cyclonic and the anticyclonic sides considered separately. Overall, dissipation asymmetry decreases with weakening rotation, approaching symmetry at $Ro\geq O(5)$ across a range of $Fr$.
In cases Fr030 and Fr040 (Fig. 9(e,f)), it is clear that the anticyclone side is strongly influenced by rotation, and that the dissipation increases rapidly as rotation strengthens until it saturates in the $Ro\leq O(0.5)$ range. On the other hand, the cyclonic side is rather weakly influenced by rotation except for cases Fr030Ro075 and Fr040Ro040. The cyclonic dissipation peaks are presumably due to the instability of the cyclonic shear layer/vortices and strong interactions between cyclonic and anticyclonic sides (see Fig 7(b) and the supplementary movies). It is also noted in Appendix C that cyclonic vortices can also become unstable to CI.
In the strongly stratified cases Fr007 and Fr015 (Fig. 9(c,d)), dissipation on the cyclonic side increases monotonically with $Ro$, suggesting that the cyclonic side is subject to the stabilizing effect of rotation. Anticyclonic dissipation, however, displays a mild peak at $Ro=0.75$ for both cases, suggesting that CI is attenuated but still leaves a signature in the dissipation. Intuitively, as stratification becomes stronger, such as in cases Fr007 and Fr015, the vertical length scale for instability decreases and the space for CI shrinks until it is eventually suppressed. The anticyclonic CI attenuation and the cyclonic stabilization add up to a net effect of rotation being solely stabilizing at $Fr=0.075,0.15$.
@@
The present parametric dependence has several implications. First, the inertial scaling $\varepsilon=C_{\varepsilon}U^{3}_{\infty}/D$, where $C_{\varepsilon}$ is a scaling coefficient, works reasonably well for the wake dissipation over a wide range in the parameter space of $(Fr,Ro)$ (the value of $FrRo$ varies by 2 orders of magnitude). It provides a means for contextualizing field measurements or modeling wake dissipation/mixing in regional or global climate models. On the other hand, the variability of $C_{\varepsilon}$ over the range of $Fr$ and $Ro$ investigated reflects systematic effects of stratification and rotation. It is evident that turbulence in seamount or hill wakes at different levels of stratification depends differently on rotation, with a transition point between $Fr=0.15$ and $Fr=0.30$ due to the activation of the CI. That being said, $C_{\varepsilon}$ is likely not a simple function of $Fr$ and $Ro$.
@@
In previous numerical studies of topographic wakes (Perfect et al., 2018, 2020; Srinivasan et al., 2021), it was suggested that the vortex dynamics and turbulent dissipation might both be categorized by a single parameter, instead of $(Fr,Ro)$. For example, Perfect et al. (2018) varied both $Fr$ and $Ro$ in the range of $0.014\leq Fr^{*}\leq 0.14$ and $0.053\leq Ro^{*}\leq 0.21$, which is equivalent to $0.014\leq Fr\leq 0.14$ and $0.0265\leq Ro\leq 0.105$ after a conversion to our definition ($Fr=Fr^{*},Ro=Ro^{*}/2$). It was suggested that the vortex dynamics can be characterized as a function of the Burger number, $Bu=(Ro/Fr)^{2}=(Nh/f_{c}D)^{2}$, and the dissipation can further be parameterized as a function of a positive power of $FrRo$. Srinivasan et al. (2021) studied topographic wakes as a function of RoRo ($0.025\leq Ro\leq 1$) when $Fr=0.02$ is fixed, and found that the dissipation monotonically increases as a function of the Rossby number. For comparison, the same data in Fig. 9(a) is plotted in Fig. 9(b), as a function of $FrRo$. Although the data seems to collapse better, the division into two groups with different dependencies is still very clear. The parameters in Perfect et al. (2020) have $(FrRo)_{\rm max}\approx 0.015$ and those in Srinivasan et al. (2021) have $(FrRo)_{\rm max}=0.02$. Both overlap with the lower end of the present parameters, as shown by the green and red shades in Fig. 9(b).
@@
It was pointed out by Liu et al. (2024) that there is still vertical coupling of the vortex shedding at $Fr=0.15,\,Bu=\infty$ and hence the $Bu$-determination of coupling/decoupling is incomplete without an additional $Fr$-dependence – stronger stratification than $Fr=0.15$ is required to vertically decouple the vortex shedding in rotating and non-rotating Kármán wakes. A similar role of stratification is found in the dissipation dependence, that the activation or attenuation of the CI between $Fr=0.15$ and $Fr=0.30$ changes the $Ro$-dependence of dissipation. Thus, $Fr$ plays a crucial role as a separate independent parameter in the regime transitions of vortex dynamics and turbulence generation. It requires separate attention instead of being simply absorbed into $Bu$ or $FrRo$. Last but not least, the asymmetry between the cyclonic and anticyclonic sides is significant at moderate to strong rotation. The difference between the responses of cyclonic and anticyclonic shear layers/vortices to changing $Ro$ requires separate consideration in the analysis and parameterization of topographic wakes.
Figure 10: Joint p.d.f (JPDF) of the total KE dissipation ($\tilde{\varepsilon}$, defined in (9)) and the local gradient Richardson number ($Ri_{g}$, defined in (10)). Contours enclose $85\%$ of the JPDF in each case. Plane location $z/h=0.25$. The FrFr-dependent black dashed lines in each figure are given by $\tilde{\varepsilon}D/U_{\infty}^{3}=C_{R}Fr^{-1}Ri_{g}^{-1}$, with the same fitting constant $C_{R}=1/2500$.
4.2 Relating dissipation rate to local stability criteria
Besides global measures such as volume-integrated dissipation as a function of $Fr$ and $Ro$, local measures such as the joint distributions of stability criteria and dissipation could provide additional information for parameterization. Possible surrogates include the gradient Richardson number for KHI and PV for CI and the overarching goal is to seek whether, given information about the local stability criteria, the prediction of dissipation can be improved.